Suppose that a firm produces at an output level where the MPL is 50 units and the wage rate is Rs 25 (w). Suppose further that the MPK is 100 units and the rental price of land is Rs.40 Q1 = Is the firm producing efficiently? Q2 = If the firm is not producing efficiently , how might it do so?


1
Expert's answer
2021-08-25T10:05:17-0400

Determinations of the forms production efficiency:

Using Least Cost Rule;


MRTS = "\frac{MPk}{MPl} = - \frac{\Delta l}{\Delta k}"


MRTS = "\frac{100}{50} = 2"

This means that the firm will have to either increase the labor (L) by 2 units and reduce the units of capital (K) by one unit, and vice versa.

If it chooses the former the costs combination (K, L) will be (-40, 50) and the latter will be (40, -50).


Conclusion; The cost combination (K,L) of (40, -50) will be appropriate for the firm since it will be able to have to save 10 on cost of inputs.